OpenStudy (anonymous):

true or false is the following a arithmetic sequence. 5, 6, 7, 8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@dg123 thanks so much!!! whats this one? What is the common difference of the following arithmetic sequence? 3, 13, 23, 33,… (sorry this is life or death right now)

OpenStudy (anonymous):

10

OpenStudy (anonymous):

omg stay online please!!!! i will thank you later!!! lol thank you

OpenStudy (anonymous):

What are the next 3 terms of the following sequence? 5.5, 6,7, 7.9, 9.1,…

OpenStudy (anonymous):

yeah its quite late in the night here in india, but i will stay online

OpenStudy (anonymous):

10.3,11.5,12.7

OpenStudy (anonymous):

What is the common difference of the following arithmetic sequence? -8, 0, 8, 16,…

OpenStudy (anonymous):

8

OpenStudy (anonymous):

What are the next 3 terms in the following sequence? 2.3, 6.6, 10.9,…

OpenStudy (anonymous):

15.2,19.5,23.8

OpenStudy (anonymous):

What is the common difference of the following arithmetic sequence? -1000, -2000, -3000,…

OpenStudy (anonymous):

-1000

OpenStudy (anonymous):

just wanna pause and say thanks again!!!! lol

OpenStudy (anonymous):

fire on

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Find the four arithmetic means between 40 and 10.

OpenStudy (anonymous):

your answer is {34,28,22,16} God bless you.

OpenStudy (anonymous):

Find the three arithmetic means between 5 and 17

OpenStudy (anonymous):

8,11,14

OpenStudy (anonymous):

What is the 5th term of an arithmetic sequence if t3 = 10 and t7 = 26?

OpenStudy (anonymous):

18

OpenStudy (anonymous):

Find the four arithmetic means between 40 and 10

OpenStudy (anonymous):

34,28,22,16}

OpenStudy (anonymous):

Find S20 of the arithmetic series 14 + 16 + 18 + 20 +…

OpenStudy (anonymous):

10(28+38)=660

OpenStudy (anonymous):

Find S10 for 10 + 8 + 6 + 4 +…

OpenStudy (anonymous):

5(20+(9*-2))=10

OpenStudy (anonymous):

Find S25 for 3 + 7 + 11 + 15 +

OpenStudy (anonymous):

25/2(6+24(4))=102*25/2=51*25=1275

OpenStudy (anonymous):

is ur exercise still left?

OpenStudy (anonymous):

yes.. sorry about this! i gotta have it done in the next hour.. i appreciate your help so much!!!!!

OpenStudy (anonymous):

Find S20 for the arithmetic series 8 + 16 + 24 + 32 +…

OpenStudy (anonymous):

10(16+19*8)=1680 how many questions left?

OpenStudy (anonymous):

20 but we can stop whenever your ready.. just let me know.

OpenStudy (anonymous):

its really d time i go to sleep or else i wont b avle to wake up early tommorrow ,sorry

OpenStudy (anonymous):

ok man thanks!!!!!! really grateful for your help. do u have twitter?? so i can follow u

OpenStudy (anonymous):

naah, facebook

OpenStudy (anonymous):

i m on facebook

OpenStudy (anonymous):

ok when get a chance add me search "james marshall baity"

OpenStudy (anonymous):

adding now:)

OpenStudy (anonymous):

ok.. get some sleep bro. thanks again

OpenStudy (anonymous):

anytime, u from chicago ?

OpenStudy (anonymous):

keep an eye for divyansh david firzen gupta, tht me, u might have a request